# DataChef: Cooking Up Optimal Data Recipes for LLM Adaptation via Reinforcement Learning

### v1

@@ 4.3 Data Verifier
To validate the proposed Data Verifier, we analyze the Pearson correlation between the verifier scores and downstream benchmark performance. We benchmark against several widely used data evaluation metrics, including IFD Li et al. (2024a), RewardModelScore Liu et al. (2025a), DEITA Liu et al. (2024), and VendiScore Friedman and Dieng (2023). To ensure diversity in data quality and model performance, we construct 8–12 datasets per task under a fixed data budget using two strategies: (1) Direct sampling from available task-specific data sources. (2) Subset selection from the pool formed in (1) based on response length. As shown in Fig. 3, with supplementary plots in Appx. D, our Data Verifier exhibits a strong positive correlation with downstream model performance across all six tasks, achieving an average Pearson correlation of $0.59$. Crucially, the Data Verifier maintains consistent positive correlation across diverse task settings, indicating superior robustness compared to baselines.
@@ A.2 Prompt Templates and Model Selection
Data Verifier. We employ gpt-oss-120b OpenAI et al. (2025) as the backbone for the Data Verifier. The detailed rubric-based prompt used for evaluation is presented below.
Cold-start Models. To construct high-quality cold-start supervision, we leverage two specialized models: Qwen3-Next-80B-A3B-Thinking Yang et al. (2025) for planning and reasoning, and Kimi-K2-Instruct Team et al. (2025) for code implementation. The specific prompts used for these roles are provided below.
Appendix B Details of Experiments Setup
B.1 Data Evaluation Metrics Settings
  1. $\bullet$
  2. $\bullet$
  3. $\bullet$
  4. $\bullet$
B.2 Evaluation Setup
  1. $\bullet$
  2. $\bullet$
  3. $\bullet$
  4. $\bullet$
  5. $\bullet$
  6. $\bullet$
@@
[…]

### v2

@@ 4.3 Data Verifier
Setup. To validate the proposed Data Verifier, we analyze the Pearson correlation between the verifier scores and downstream benchmark performance. We benchmark against several widely used data evaluation metrics, including IFD Li et al. (2024a), RewardModelScore Liu et al. (2025a), DEITA Liu et al. (2024), and VendiScore Friedman and Dieng (2023). To ensure diversity in data quality and model performance, we construct 8–12 datasets per task under a fixed data budget using two strategies: (1) Direct sampling from available task-specific data sources. (2) Subset selection from the pool formed in (1) based on response length.
Correlation Analysis. As shown in Fig. 3 and detailed in Appx. D, the Data Verifier exhibits superior capability compared to existing metrics, achieving the highest average Pearson correlation of $0.59$ across all six domains. Crucially, while baseline metrics suffer from severe cross-domain variance, frequently yielding negative correlations that can provide misleading optimization signals in domains such as Math (0.48-0.48 for IFD) and Code (0.15-0.15 for DEITA), our Data Verifier maintains a strictly positive correlation across all evaluated tasks. Furthermore, it demonstrates the highest statistical significance, with 4 out of 6 tasks satisfying $p<0.1$, confirming that the Data Verifier provides a statistically robust and globally consistent signal for automated data recipe generation.
@@ A.1 Details of Task Pool
Full Benchmark List. Table 4 presents the full list of benchmarks used in our task pool, along with their domains and usage.
@@
Datasets Retrieval Procedure. We automate the retrieval of candidate datasets relevant to the benchmark and task through the following procedure:
  1. $\bullet$
    Keyword Synthesis. We use an LLM to generate 3–5 high-relevance search keywords tailored to the downstream task.
  2. $\bullet$
    Search. Leveraging the generated keywords, we query both the Hugging Face and Google Search APIs to harvest a broad spectrum of potential datasets.
  3. $\bullet$
    Rank. To ensure dataset utility, we rank retrieved datasets by community popularity (e.g., likes) and retain the top-4 candidates for each keyword.
  4. $\bullet$
    Verify. We implement a strict verification protocol to preclude data leakage, explicitly ensuring no overlap exists between the retrieved candidates and the benchmark.
@@ A.2 Prompt Templates and Model Selection
[…]
Figure 7: Prompt used for Data Verifier.
Data Verifier. We utilize gpt-oss-120b OpenAI et al. (2025) as the Data Verifier. This is a Mixture-of-Experts (MoE) model with only about 5B active parameters per token, ensuring high inference speed. The detailed rubric-based prompt used for evaluation is presented in Fig. 7.
@@
Figure 8: Prompt used for data recipe generation (Natural Language).
@@
Figure 9: Prompt used for data recipe generation (Executable Code).
Cold-start Models. To construct high-quality cold-start supervision, we employ two specialized models: Qwen3-Next-80B-A3B-Thinking Yang et al. (2025) for planning and reasoning, and Kimi-K2-Instruct Team et al. (2025) for code implementation. The corresponding prompts are detailed in Fig. 8 and Fig. 9.
Appendix B Details of Experiments Setup
B.1 Data Evaluation Metrics Settings
  1. $\bullet$
  2. $\bullet$
  3. $\bullet$
  4. $\bullet$
Table 5: Detailed correlation results across different domains. For each domain, we report the Pearson correlation coefficient ($r$) and the corresponding $p$-value.
Metric PHYSICS AIME LiveCodeBench ClimaQA OpenFinData CHID Summary
$r$ $p$ $r$ $p$ $r$ $p$ $r$ $p$ $r$ $p$ $r$ $p$ Positive $p<0.05$ $p<0.1$
IFD 0.48 0.19 -0.48 0.12 0.43 0.25 0.90 0.00 0.59 0.10 0.51 0.16 5/6 1/6 2/6
Deita 0.20 0.61 -0.04 0.91 -0.15 0.70 0.91 0.00 0.10 0.80 0.61 0.08 4/6 1/6 2/6
RewardModel 0.34 0.37 0.18 0.58 0.30 0.43 0.90 0.00 0.36 0.34 0.44 0.23 6/6 1/6 1/6
VendiScore 0.36 0.34 0.33 0.29 -0.17 0.67 0.66 0.08 0.60 0.09 0.29 0.44 5/6 0/6 2/6
DataVerifier 0.63 0.07 0.59 0.04 0.63 0.07 0.52 0.19 0.34 0.37 0.83 0.01 6/6 2/6 4/6
B.2 Evaluation Setup
  1. $\bullet$
  2. $\bullet$
  3. $\bullet$
  4. $\bullet$
  5. $\bullet$
  6. $\bullet$
B.3 Oracle Selection
The specific procedure for Oracle Selection is:
  1. 1. Filter: From the 32 generated recipes, we select the top-8 candidates based on the Data Verifier Score.
  2. 2. Verify: A human expert reviews these top-8 candidates to ensure quality, verifying key aspects including: (a) format alignment with the target benchmark (e.g., the processed dataset should include multiple-choice structures for MCQ tasks); (b) context integrity (e.g., verifying that inputs include necessary reference texts for reading comprehension, rather than just questions); and (c) logical comprehensiveness of the pipeline operations. Based on this comprehensive assessment, the expert selects the optimal candidate for final training.
B.4 Budget
For DBS evaluation, the data budget is set to 10K. If a recipe generates more, we randomly downsample to 10K; if fewer, we use all. During training and DVS evaluation, we use a budget of 100 samples for rapid proxy reward calculation.
@@ Appendix D Additional Results on Correlation Analysis
Appendix E Computational Cost Analysis
For SFT, the average per-step time is 2.4 minutes with a batch size of 32 on 8 H200 GPUs. For RL, we allocate 8 H200 GPUs to the policy model and 2 H200 GPUs to the Data Verifier deployment. A single RL step, which processes 128 candidate recipes, takes an average of 20.2 minutes. Notably, the Data Verifier inference accounts for less than 2 minutes of this duration, demonstrating its high computational efficiency.